# MathEDU: Towards Adaptive Feedback for Student Mathematical Problem-Solving

### v1

@@ 6.1 Experimental Setup
We examined the performances of Llama3 8B (Touvron et al., 2023), Llama3 70B, and GPT-3.5 with few-shot prompting. [^9] The version we used is gpt-3.5-turbo-0125. We further fine-tuned Llama3 8B with a learning rate set to 2e-4 and a LoRA rank of 16.101010All the prompts are shown in Appendix E. During inference, the temperature was set to 0, with a maximum output of 512 tokens, while other parameters remained at their default settings.
Data Splitting. To evaluate the model’s performance across different application scenarios, we designed two dataset splitting settings. (1) The first setting employs a time-series split, simulating a situation where the model has access to a student’s prior answer history. Student records were organized chronologically, with part of the records serving as the training set (answer history) and the remaining records as the validation and test sets in a 70:15:15 ratio. This approach assesses whether the model can leverage past responses to identify a student’s strengths and weaknesses and apply this knowledge to grade new answers. (2) The second setting also uses a time-series split but applies a leave-one-out strategy to simulate the model encountering a new student without prior history. In this setup, one student’s data is designated as the test set, while the data from the remaining students is proportionally divided into training and validation sets in a 9:1 ratio. This scenario, akin to the cold-start setting, evaluates the model’s adaptability to diverse problem-solving styles. [^11] The data splitting statistics are in Appendix F.
Inference Prompt. For the few-shot prompting, six examples were provided, consisting of three correct and three incorrect student responses. In the time-series split setting, we randomly selected the student’s prior answer records from the training set as few-shot demonstrations. In the leave-one-out setting, few-shot demonstrations were randomly sampled from the answer records of other students. For the fine-tuned model, which was extensively trained on the task, we employed zero-shot prompting to generate results directly.
Time Series Split Leave-One-Out
All General Gain Physics Geo. Prob. Other All General Gain Physics Geo. Prob. Other
Llama3 8B w/o $r$ 62.02% 64.66% 64.86% 55.56% 77.50% 62.50% 48.48% 63.53% 66.57% 64.61% 58.50% 69.21% 55.06% 50.31%
w/ $r$ 70.15% 71.55% 68.47% 68.42% 80.00% 75.00% 60.61% 68.58% 71.59% 68.02% 63.81% 75.47% 61.31% 61.45%
Llama3 70B w/o $r$ 83.58% 82.76% 88.29% 79.53% 97.50% 56.25% 90.91% 85.25% 86.54% 87.90% 80.89% 87.53% 74.70% 84.92%
w/ $r$ 81.76% 78.88% 87.39% 79.53% 97.50% 62.50% 84.85% 84.12% 84.57% 87.92% 81.01% 82.66% 63.69% 85.29%
GPT-3.5 w/o $r$ 40.96% 40.95% 45.05% 38.01% 52.50% 50.00% 24.24% 41.70% 43.45% 46.03% 36.91% 41.67% 37.50% 32.76%
w/ $r$ 50.41% 50.00% 54.95% 49.12% 57.50% 43.75% 39.39% 49.10% 51.65% 52.90% 43.61% 48.25% 59.23% 37.68%
Single-task w/o $r$ 87.23% 85.77% 90.09% 88.30% 90.00% 68.75% 87.88% 71.93% 71.43% 72.00% 71.98% 72.51% 72.32% 73.62%
w/ $r$ 92.70% 90.09% 95.50% 92.98% 95.00% 87.50% 100.00% 92.45% 92.45% 93.32% 91.73% 92.42% 88.99% 93.57%
Multi-task w/o $r$ 73.13% 70.26% 71.17% 80.70% 67.50% 56.25% 75.76% 43.07% 42.41% 41.95% 45.03% 44.03% 40.48% 41.91%
w/ $r$ 78.11% 74.57% 80.18% 83.63% 75.00% 62.50% 78.79% 54.77% 53.71% 53.87% 55.78% 58.00% 64.88% 55.99%
End-to-end w/o $r$ 87.73% 84.05% 93.69% 90.06% 90.00% 62.50% 90.91% 82.91% 82.29% 83.80% 83.67% 83.42% 75.00% 80.15%
w/ $r$ 91.54% 88.79% 95.50% 93.57% 95.00% 62.50% 96.97% 93.13% 93.01% 93.31% 92.53% 93.97% 95.83% 94.61%
Table 4: Results of Answer Accuracy Assessment.
@@ 6.2 Experimental Results
Answer Accuracy Assessment. We evaluated the ability of LLMs to determine the correctness of student answers. Accuracy is adopted as the evaluation metric. The results are shown in Table 4. We tested whether the detailed rationales in MathQA assist LLMs in grading student answers. “w/o $r$” and “w/ $r$” refer to whether the input includes the rationale of each question, respectively. In the time-series split setting, Llama3 70B achieved 83.58% accuracy, significantly outperforming Llama3 8B (62.02%) by 21.56%, highlighting the importance of parameter size in mathematical computation and understanding. GPT-3.5’s lower accuracy stemmed from its strict evaluation of omitted calculations. For example, if a student’s final answer was correct but some intermediate steps were skipped, GPT-3.5 often classified the answer as incorrect.
@@
The bottom three rows of Table 4 represent our fine-tuned models, which demonstrate higher accuracy in determining correctness compared to the original, non-fine-tuned Llama3 8B model. Among these, the end-to-end training method perform better than the multi-task training method. This improvement may be attributed to the end-to-end approach, which requires the model to generate grading results sequentially, enabling it to perform more coherent reasoning and achieve better outcomes.
The impact of including rationales in the input is evident in the results shown in Table 4. For instance, Llama3 8B’s accuracy increased from 62.02% to 70.15%. Notably, Llama3 70B exhibited a slight decrease in accuracy from 83.58% to 81.76%. In general, the inclusion of rationales allowed most models to better assess correctness by providing additional contextual support, though the ability to interpret and utilize these rationales varied across models. Fine-tuned models, particularly the end-to-end model, demonstrated a stronger ability to effectively leverage rationales
In the leave-one-out setting, non-fine-tuned LLMs achieved accuracy consistent with the time-series split. In contrast, fine-tuned multi-task models saw a 20% accuracy drop and struggled with unseen problem-solving approaches, occasionally failing to produce outputs. However, the end-to-end model remained stable, achieving the highest accuracy of 93.13%, adapting effectively to solving styles excluded from training. We further conducted McNemar’s test to assess the significance of differences between our fine-tuned models and the baseline Llama3 70B without rationale. In the time-series split setting, the single-task model with rationale significantly outperformed the baseline, with $p<0.001$. Similarly, in the leave-one-out setting, the end-to-end model with rationale also achieved significantly better performance than the baseline, with $p<0.001$.
Although the end-to-end model seems to perform worse in the time-series split setting than in the leave-one-out setting, this likely reflects differences in data distributions, making direct comparison invalid. Moreover, the lack of user embeddings and effective encoding of individual records may limit the model’s ability to capture answering patterns, which we leave as future work.
Time Series Split Leave-One-Out
EM ($\uparrow$) Distance ($\downarrow$) EM ($\uparrow$) Distance ($\downarrow$)
Llama3 8B w/o $r$ 28.40% 68.44 10.58% 96.56
w/ $r$ 11.73% 97.73 8.80% 102.12
Llama3 70B w/o $r$ 32.10% 67.35 23.30% 78.33
w/ $r$ 29.63% 73.38 24.34% 78.55
GPT-3.5 w/o $r$ 30.25% 57.93 27.81% 63.07
w/ $r$ 27.78% 62.32 29.46% 63.26
Single-task w/o $r$ 24.69% 95.64 23.41% 96.85
w/ $r$ 23.46% 97.21 21.98% 99.09
Multi-task w/o $r$ 23.46% 95.70 22.59% 93.77
w/ $r$ 17.28% 96.35 22.11% 97.92
End-to-end w/o $r$ 30.25% 71.48 29.39% 74.01
w/ $r$ 40.12% 51.78 42.81% 52.21
Table 5: Results of Problem-Solving Error Identification.
Time Series Split Leave-One-Out
ROUGE-1 ROUGE-2 ROUGE-L BERTScore LLM Rating ROUGE-1 ROUGE-2 ROUGE-L BERTScore LLM Rating
Llama3 8B w/o $r$ 0.2593 0.0809 0.2391 0.7654 2.68 0.2103 0.0615 0.1924 0.6851 2.38
w/ $r$ 0.1819 0.0557 0.1632 0.5515 2.14 0.1738 0.0515 0.1594 0.5600 1.97
Llama3 70B w/o $r$ 0.2587 0.0881 0.2439 0.7147 3.61 0.2412 0.0761 0.2254 0.7219 3.77
w/ $r$ 0.2456 0.0876 0.2311 0.7108 3.31 0.2367 0.0720 0.2201 0.7205 3.56
GPT-3.5 w/o $r$ 0.2897 0.1041 0.2695 0.8395 3.70 0.2824 0.0978 0.2601 0.8477 3.39
w/ $r$ 0.2848 0.1025 0.2597 0.8366 3.63 0.2759 0.0925 0.2533 0.8451 3.31
Single-task w/o $r$ 0.1661 0.0730 0.1594 0.4836 0.63 0.0886 0.0359 0.0841 0.2639 0.47
w/ $r$ 0.0790 0.0378 0.0750 0.2182 0.39 0.0594 0.0205 0.0561 0.1828 0.33
Multi-task w/o $r$ 0.0130 0.0049 0.0124 0.0437 0.08 0.0600 0.0237 0.0577 0.1832 0.31
w/ $r$ 0.1650 0.0719 0.1552 0.4372 0.55 0.0244 0.0068 0.0230 0.0807 0.22
End-to-end w/o $r$ 0.1862 0.0751 0.1761 0.5327 0.99 0.1808 0.0637 0.1720 0.5726 2.14
w/ $r$ 0.2842 0.1418 0.2703 0.6786 1.91 0.2779 0.1089 0.2614 0.7940 2.28
@@
Problem-Solving Error Identification. To evaluate the model’s performance in identifying erroneous equations in cases where a student’s answer is incorrect, we employ two metrics. The first is the exact match (EM) ratio, which measures the correspondence between model-detected step and manually annotated step. The second metric is the distance, defined as the textual distance between the model-detected errors and the manually labeled errors. Textual distance measures the character difference between the predicted and annotated errors, reflecting the model’s prediction deviation. [^12] The details of the distance metric is in Appendix G.
The results are shown in Table 5. In the time series split setting, non-fine-tuned LLMs achieved an EM ratio of around 30% without rationale. However, providing rationale reduced the EM ratio for all models, with Llama3 8B showing the largest drop of 16.67%. This may result from differing formats between the rationale and the student’s actual process (e.g., -formatted), which increases the complexity of error identification.
For fine-tuned models, the single-task model performed the worst without rationale. It rarely identified errors correctly, often gave irrelevant responses, and sometimes failed to output answers. Its performance further declined when rationale was included. The multi-task model showed a similar pattern. In contrast, the end-to-end model outperformed Llama3 8B without rationale and showed noticeable improvements with rationale. This result indicates that fine-tuning with this approach helps the model better understand and utilize rationale for error identification.
In the leave-one-out setting, Llama3 8B with few-shot prompting performed poorly. Other LLMs performed similarly to their results in the time-series split method. This result indicates the limitation of smaller parameter models in adapting to different scenarios. Among fine-tuned models, the end-to-end model performed consistently, adapting well to unseen question types and problem-solving styles.
Feedback Generation. We evaluate the quality of these suggestions using five metrics: ROUGE-1, ROUGE-2, ROUGE-L (Lin, 2004), BERTScore (Zhang et al., 2019), and LLM Rating (Liu et al., 2023). The LLM Rating, using GPT-4 as the evaluator, compares the model generated feedback to the ground truth and assigns a score ranging from 0 to 5, with higher scores indicating better quality. The details of the scoring criteria are in Appendix G. The results in Table 6 show that non-fine-tuned language models perform consistently across the time series and leave-one-out splits. Their inability to effectively utilize additional rationales for math problems limits their ability to generate accurate suggestions.
@@
In our fine-tuned models, the single-task and multi-task models performed poorly, rarely identifying errors or generating feedback. Although the end-to-end models achieved ROUGE and BERTScore results comparable to Llama3 70B, they received lower LLM ratings. This represents the complexity of feedback generation, which requires larger models with stronger reasoning and language generation capabilities. Even with fine-tuning, Llama3 8B appears to have reached its performance limit. We also explored the capabilities of o1-mini but conducted only partial experiments due to high financial costs. Main experiments for each task were completed, excluding the leave-one-out and without rationale settings. Results and discussion are in Appendix H.
7 Discussion
7.1 Relationship Between Problem-Solving Ability and Grading Ability
One concern about using LLMs in the math tutoring system is that they may act differently in problem-solving and grading students’ answers. To address this issue, we investigate the correlation between LLMs’ mathematical problem-solving abilities and their accuracy in grading students’ processes. [^13] The model’s performance is presented in Appendix I. A Chi-square test was conducted using predictions from the time-series split setting, excluding rationales to rely solely on the model’s intrinsic capabilities. Since the Chi-square test requires categorical variables, we defined task-specific criteria for statistical validity. Answer accuracy assessment is a binary classification task, allowing direct Chi-square application. Error identification is an extraction task where only exact matches with the ground truth are correct; all others are incorrect, ensuring the results are treated as categorical variables. Feedback generation lacks an absolute correctness measure, so we computed ROUGE-L scores for all generated feedback and set the average score as the threshold, considering feedback correct if it exceeded this value. We report the corresponding $p$-values in Table 7.
Answer Acc. Error Ident. Feedback Gen.
Llama3 8B 1.0000 0.2191 0.8819
Llama3 70B 0.4401 0.3414 0.7407
GPT-3.5 0.8410 0.8538 0.8640
End-to-end 0.3766 0.8852 0.5567
Table 7: Correlation between LLM’s Math Word Problem Solving Ability and Grading Ability.
The results show no significant correlation between the model’s problem-solving ability and its performance in grading tasks. Grading requires more than aligning solutions; it involves mathematical reasoning and step-by-step analysis to assess each part of the student’s process. Effective feedback generation further demands understanding the error, its cause, and providing accurate, context-specific corrections. These findings emphasize the necessity of advanced reasoning and a thorough understanding of mathematical concepts. We also analyzed whether different problem-solving styles among students affect the model’s grading performance. The results show that LLMs perform better when solutions include more textual explanations. Detailed discussions are provided in Appendix J.
7.2 Types of Student Errors that Models Excel at Resolving
To further evaluate the model responses, we selected representative results from Llama3 70B without rationale, GPT-3.5 without rationale, and an end-to-end model with rationale under the time series split setting. Additionally, we included the o1-mini model, which uses few-shot prompts without rationale, to further enhance the comparison and analysis of model performance. The invited three mathematics experts were asked to evaluate the quality of the generated feedback. The evaluation criteria range from 0 (not helpful) to 3 (clear and effective guidance). [^14] The evaluation criteria is elaborated in Appendix K.
@@ 7.2 Types of Student Errors that Models Excel at Resolving
Table 8: Human Evaluation of Generated Feedback from Different Models. “Wrong M.,” “Calc.,” “Incomp.,” “Lack of M.,” and “Careless,” correspond to “Wrong Mathematical Operation/Concept,” “Calculation Error,” “Incomplete Answer,” “Lack of Necessary Mathematical Concepts,” and “Careless Error,” respectively.
The results are shown in Table 8.151515The examples and the corresponding analyses are presented in Appendix L. We assessed the inter-rater agreement on the models’ output ratings using Krippendorff’s Alpha, yielding a score of 0.7628, which reflects a moderate level of agreement among the experts. The responses from language models remain below the ideal score of 3. The fine-tuned end-to-end model underperformed compared to larger models. All models excelled in addressing “Lack of Necessary Mathematical Concept” errors. The reason may be that such errors arise when students cannot solve the problem, requiring the model to solve it directly and explain, without analyzing the student’s process.
All models struggled with identifying “Calculation Errors,” often failing even with simple arithmetic mistakes. Interestingly, they performed better on “Wrong Mathematical Operations/Concepts.” Models tended to assume student results were correct and frequently misattributed errors to incorrect methods. In some cases, they even revised correct processes, producing inaccurate feedback. Our fine-tuned end-to-end model performed worse than larger models in “Wrong Mathematical Operations/Concepts,” “Incomplete,” and “Careless Errors,” likely due to its limited reasoning and process analysis capabilities. This led to inaccurate judgments on errors requiring deeper understanding. However, the performance gap narrowed in categories demanding less reasoning, such as “Calculation Errors” and “Lack of Necessary Mathematical Concepts.”
Although GPT-3.5, o1-mini, and Llama3 70B have stronger reasoning and generation capabilities, they struggle to identify student errors and provide focused feedback. Larger models, like o1-mini, often produce verbose responses with inaccuracies, which experts find overwhelming for students and ineffective in addressing specific mistakes. In contrast, our fine-tuned 8B end-to-end model provides concise feedback that directly targets errors but performs poorly on unfamiliar problem types outside its training data. Large models (e.g., GPT-3.5) also tend to solve problems themselves rather than offering adaptive feedback tailored to the student’s reasoning process, even when prompts emphasize error analysis. These results highlight the need for further refinement in LLMs, particularly in handling more nuanced aspects of students’ problem-solving processes.

### v2

@@ 5.1 Experimental Setup
Models. We evaluated Llama3 8B (Touvron et al., 2023), Llama3 70B, and GPT-3.5 (Ouyang et al., 2022) [^13] The version we used is gpt-3.5-turbo-0125. using few-shot prompting. We also explored the capabilities of the reasoning model (i.e., o1-mini (Jaech et al., 2024)). We further fine-tuned Llama3 8B with a learning rate set to 2e-4 and a LoRA rank of 16.141414All the prompts are shown at: https://github.com/NYCU-NLP-Lab/MathEDU/blob/main/Prompts.pdf. During inference, the temperature was set to 0, with a maximum output of 512 tokens, while other parameters remained at their default settings.
Data Splitting. Student records were organized chronologically, with part of the records serving as the training set (answer history) and the remaining records as the validation and test sets in a 70:15:15 ratio. This setting assesses whether the model can leverage past responses to identify a student’s strengths and weaknesses and apply this knowledge to grade new answers. Specifically, the training, validation, and test sets consist of 2,836, 609, and 603 examples, respectively.
Inference Prompt. For the few-shot prompting, six examples were provided, consisting of three correct and three incorrect student responses. We randomly selected the student’s prior answer records from the training set as few-shot demonstrations. For the fine-tuned model, which was extensively trained on the task, we employed zero-shot prompting to generate results directly.
@@ 5.2.1 Answer Correctness Classification
We evaluated the ability of LLMs to determine the correctness of student answers. F-score is adopted as the evaluation metric to address class imbalance, as correct answers are more frequent than incorrect ones. The results are shown in Table 3.151515Note that the results of the two tasks are combined in a single table. For answer correctness classification, “All” indicates the overall performance. We tested whether the detailed rationales in MathQA assist LLMs in grading student answers. “w/o rr” and “w/ rr” refer to whether the input includes the rationale of each question, respectively. Llama3 70B achieved 88.23% F-score, significantly outperforming Llama3 8B (66.96%), indicating the importance of parameter size in mathematical computation and understanding. GPT-3.5’s lower F-score stemmed from its strict evaluation of omitted calculations. For example, if a student’s final answer was correct but some intermediate steps were skipped, GPT-3.5 often classified the answer as incorrect. Notably, o1-mini demonstrates exceptional mathematical capabilities, achieving an overall F-score of 94.66% in identifying the correctness of the student’s answers. This represents the value of strong reasoning capabilities in correctness identification. [^16] We did not conduct experiments with rationales, as they substantially increased input length, led to longer outputs, and reduced output stability.
@@
The bottom three rows of Table 3 represent our fine-tuned models, which demonstrate higher F-score in determining correctness compared to the original, non-fine-tuned Llama3 8B model. Among these, the end-to-end training method perform better than the multi-task training method. This improvement may be attributed to the end-to-end approach, which requires the model to generate grading results sequentially, enabling it to perform more coherent reasoning and achieve better outcomes.
The impact of including rationales in the input is evident in the results. For example, Llama3 8B’s F-score rose from 66.96% to 77.04%, while Llama3 70B showed a slight decline. Fine-tuned models demonstrated a stronger ability to effectively leverage rationales. We further conducted McNemar’s test to assess the significance of differences between our fine-tuned models and the baseline Llama3 70B without rationale. The single-task model with rationale significantly outperformed the baseline, with $p<0.001$. Overall, leveraging LRMs can yield strong performance on this task, but for scenarios requiring a smaller-scale, more controllable system, fine-tuning a small LLM offers a cost-effective alternative that can even surpass the performance of much larger models.
5.2.2 Problem-Solving Error Identification
To evaluate the model’s performance in identifying erroneous equations in cases where a student’s answer is incorrect, we employ two metrics. The first is the exact match (EM) ratio, which measures the correspondence between model-detected step and manually annotated step. The second metric is “Distance,” computed using the Hausdorff distance between the model-detected and manually annotated error spans. This metric captures the deviation between predicted and gold spans, even when they do not overlap exactly.
The results are shown in Table 3. Non-fine-tuned LLMs achieved an EM ratio of around 30% without rationale. However, providing rationale reduced the EM ratio for all models, with Llama3 8B showing the largest drop of 16.67%. This may result from differing formats between the rationale and the student’s actual process (e.g., -formatted), which increases the complexity of error identification. For fine-tuned models, the multi-task model performed the worst without rationale. It rarely identified errors correctly, often gave irrelevant responses, and sometimes failed to output answers. Its performance further declined when rationale was included. The single-task model showed a similar pattern. In contrast, the end-to-end model outperformed Llama3 8B without rationale and showed noticeable improvements with rationale. This result indicates that fine-tuning with this approach helps the model better understand and utilize rationale for error identification.
o1-mini demonstrated stronger mathematical reasoning and understanding abilities, surpassing the end-to-end model. However, it performed worse in terms of the distance metric. Upon reviewing o1-mini’s outputs, we found that in cases where errors were due to unfinished problems, the student’s calculations were correct. Teachers typically did not label any formulas as incorrect. Despite this, o1-mini still misidentified incorrect formulas in these cases, which introduced significant penalties in the distance metric, leading to a poorer performance in this aspect. This is a critical risk in educational settings, since over-diagnosing errors may unfairly penalize students and decrease trust in the system, and also reduce their willingness to learn.
ROUGE-L BERTScore LLM Rating
Llama3 8B w/o $r$ 0.2391 0.7654 2.68
w/ $r$ 0.1632 0.5515 2.14
Llama3 70B w/o $r$ 0.2439 0.7147 3.61
w/ $r$ 0.2311 0.7108 3.31
GPT-3.5 w/o $r$ 0.2695 0.8395 3.70
w/ $r$ 0.2597 0.8366 3.63
o1-mini w/o $r$ 0.1839 0.7986 4.70
Single-task w/o $r$ 0.1594 0.4836 0.63
w/ $r$ 0.0750 0.2182 0.39
Multi-task w/o $r$ 0.0124 0.0437 0.08
w/ $r$ 0.1552 0.4372 0.55
End-to-end w/o $r$ 0.1761 0.5327 0.99
w/ $r$ 0.2703 0.6786 1.91
@@ 5.2.2 Problem-Solving Error Identification
5.2.3 Feedback Generation
We evaluate the quality of these suggestions using three metrics: ROUGE-L  (Lin, 2004), BERTScore (Zhang et al., 2019), and LLM Rating (Liu et al., 2023). The LLM Rating, using GPT-4 as the evaluator, compares model-generated feedback against the ground-truth teacher suggestions and assigns a score from 0 to 5, where higher scores reflect better quality. Following established evaluation practices, GPT-4 is instructed to assess each response based on its relevance, helpfulness, and alignment with teacher feedback. A score of 0 indicates completely irrelevant or nonsensical suggestions, while a score of 5 reflects highly relevant, helpful, and well-aligned guidance. Table 4 presents the results of feedback generation. [^17] Additional results for ROUGE-1 and ROUGE-2 are provided in Appendix E. Larger models such as Llama3 70B and GPT-3.5 do not outperform Llama3 8B as expected, suggesting that generating teacher-like feedback remains a challenging task even for high-capacity models.
@@ 5.2.3 Feedback Generation
In our fine-tuned models, the single-task and multi-task models performed poorly, rarely identifying errors or generating feedback. Although the end-to-end model with rationale achieved ROUGE and BERTScore results comparable to Llama3 70B, they received lower LLM ratings. This represents the complexity of feedback generation, which requires larger models with stronger reasoning and language generation capabilities. Even with fine-tuning, Llama3 8B appears to have reached its performance limit. Specifically, the fine-tuned model excels at generating feedback for problem types it has been trained on, providing accurate suggestions. However, when confronted with unfamiliar problem types, its outputs tend to follow learned patterns, which may result in feedback that is less relevant or less accurate.
In the LLM rating conducted by using GPT-4, o1-mini received an almost perfect score of 4.70. However, its responses were often overly lengthy, generating additional irrelevant information, which impacted the evaluation results and further reduced their effectiveness in addressing specific student errors. Further discussion will be provided in the following section.
@@
Table 5: Human Evaluation of Generated Feedback Across Error Categories.
Figure 4: Feedback Categories by Models
6 Discussion
6.1 Human Evaluation
To evaluate the model responses, we selected representative results from Llama3 70B without rationale, GPT-3.5 without rationale, and an end-to-end model with rationale. The o1-mini model, which uses few-shot prompts without rationale, is included. The invited three mathematics experts were asked to evaluate the quality of the generated feedback. The evaluation criteria range from 0 (not helpful) to 3 (clear and effective guidance).  [^18] The evaluation criteria is elaborated in Appendix F.
@@ 6.1 Human Evaluation
The results are shown in Table 5. “Wrong M.,” “Calc.,” “Incomp.,” “Lack of M.,” and “Careless,” correspond to “Wrong Mathematical Operation/Concept,” “Calculation Error,” “Incomplete Answer,” “Lack of Necessary Mathematical Concepts,” and “Careless Error,” respectively. We assessed the inter-rater agreement on the models’ output ratings using Krippendorff’s Alpha, yielding a score of 0.7628, which reflects a moderate level of agreement among the experts.
The responses from language models remain below the ideal score of 3. All models excelled at handling “Lack of Necessary Mathematical Concept” errors, since these cases only require generating the correct solution and explanation. By contrast, they struggled with “Calculation Errors,” often missing even simple mistakes, and sometimes misattributing errors or revising correct processes. Our end-to-end model further lagged behind larger models on “Wrong Operations/Concepts,” “Incomplete,” and “Careless Errors,” reflecting its limited reasoning and process analysis capabilities.
Although GPT-3.5, o1-mini, and Llama3 70B have stronger reasoning and generation capabilities, they struggle to identify student errors and provide focused feedback. o1-mini often generates lengthy but inaccurate responses, while GPT-3.5 tends to solve problems directly rather than analyze student reasoning. In contrast, our end-to-end model offers concise, targeted feedback but struggles with unfamiliar problem types. These findings highlight the need for refining LLMs to better address the nuances of student problem-solving processes. Besides, experts noted that Llama3 70B occasionally produced feedback with a judgmental tone, such as “The student doesn’t know how to set up the equation.” Such phrasing, while factually descriptive, risks discouraging students rather than guiding improvement, underscoring the importance of pedagogical sensitivity in feedback generation. Additional case studies of feedback generation are provided in Appendix G.
Model S1 S2 S3 S4 S5 S6
Llama3 8B w/o $r$ 59.59% 70.36% 61.50% 67.88% 66.72% 55.15%
Llama3 8B w/ $r$ 62.37% 72.26% 64.90% 71.06% 73.90% 66.97%
End-to-end w/o $r$ 87.12% 57.96% 85.55% 87.73% 90.03% 89.09%
End-to-end w/ $r$ 94.58% 95.33% 90.86% 89.85% 96.48% 91.67%
Average 75.92% 73.98% 75.70% 79.13% 81.78% 75.72%
Table 6: Accuracy of Correctness Classification Across Students (S1–S6).
Model S1 S2 S3 S4 S5 S6
EM ($\uparrow$) Dis. ($\downarrow$) EM ($\uparrow$) Dis. ($\downarrow$) EM ($\uparrow$) Dis. ($\downarrow$) EM ($\uparrow$) Dis. ($\downarrow$) EM ($\uparrow$) Dis. ($\downarrow$) EM ($\uparrow$) Dis. ($\downarrow$)
Llama3 8B w/o $r$ 15.42% 48.69 3.53% 174.22 7.65% 137.87 4.91% 140.47 13.33% 66.20 18.60% 37.30
Llama3 8B w/ $r$ 11.44% 50.36 8.24% 187.19 4.59% 136.57 0.61% 150.21 9.33% 73.39 18.60% 41.12
End-to-end w/o $r$ 38.81% 40.51 2.35% 199.51 20.92% 117.27 20.25% 90.90 41.33% 46.28 52.71% 27.14
End-to-end w/ $r$ 58.71% 23.53 22.35% 96.47 28.57% 97.27 30.06% 92.64 61.33% 20.88 55.81% 24.50
Average 31.10% 40.77 9.12% 164.35 15.43% 122.25 13.96% 118.55 31.33% 51.69 36.43% 32.51
Table 7: Model Performance on Error Identification Across Students (Exact Match and Hausdorff Distance).
Model S1 S2 S3 S4 S5 S6
R-L Rating R-L Rating R-L Rating R-L Rating R-L Rating R-L Rating
Llama3 8B w/o $r$ 0.2269 2.98 0.1967 2.95 0.1972 2.99 0.1689 2.87 0.1945 2.35 0.1701 2.14
Llama3 8B w/ $r$ 0.1786 3.00 0.1459 2.86 0.1807 3.06 0.1372 2.67 0.1655 2.55 0.1487 1.97
End-to-end w/o $r$ 0.1738 2.16 0.0425 1.68 0.1615 2.08 0.2165 2.64 0.2040 1.66 0.2339 1.51
End-to-end w/ $r$ 0.2885 2.60 0.2676 2.51 0.2557 2.63 0.2039 3.00 0.3226 2.67 0.2301 2.00
Average 0.2169 2.69 0.1632 2.50 0.1988 2.69 0.1816 2.80 0.2216 2.31 0.1957 1.91
Table 8: Model Performance on Feedback Generation Across Students (ROUGE-L and LLM Rating).
6.2 Generated Feedback Characterization
Following Section 6.1, we extended the analysis of feedback generation with an automated, LLM-assisted approach to better capture pedagogical risks. Given the scale of samples and models, fully manual annotation was impractical. Inspired by Chen et al. (2025), we employed GPT-4o [^19] The version used is gpt-4o-2024-11-20. as a judge, evaluating each instance five times and consolidating results to ensure stability. Finally, all final results were manually reviewed. We categorized model-generated feedback errors into the following types: (1) Aligned: Feedback closely matches teacher-written feedback. (2) Redundant detail: Includes excessive, non-essential content that reduces focus. (3) Imprecise wording: Uses inaccurate or inconsistent terminology (e.g., confusing “area” with “volume”), creating ambiguity for students. (4) Missing steps: Omits one or more essential reasoning steps. (5) Calculation error: Produces incorrect numerical results. (6) Misjudged answer: Incorrectly classifies the student’s final answer as right or wrong. (7) Misidentified error: Correct feedback but pointing to the wrong error location. (8) Misunderstood approach: Misinterprets the student’s problem-solving strategy.
The results are shown in Figure 4. Overall, alignment with teacher feedback was rare; o1-mini often produced lengthy but less error-prone responses; GPT-3.5 and Llama3 70B often skipped steps or made calculation mistakes; our end-to-end model struggled with precise wording and misjudgments, indicating weaker generalization. All models exhibited reliability issues, such as misidentifying errors or misunderstanding student reasoning.
7 Impact of Student Proficiency on Model Performance
To address the variability across individual students, we report model performance on answer correctness classification, error identification, and feedback generation for each of the six students in Tables 6, 7, and 8, respectively. We focus on Llama3 8B and our end-to-end fine-tuned model, as they represent the baseline few-shot prompting approach and our best-performing fine-tuning strategy, allowing for a direct comparison between these two paradigms. Based on overall correctness rates reported in Table 16, we categorize students into three proficiency levels: high-performing (S2: 87.59%, S6: 80.61%), medium-performing (S4: 75.30%, S3: 71.09%, S1: 70.57%), and lower-performing (S5: 67.01%).
Model performance does not exhibit a clear correlation with student proficiency levels. S2, despite being the highest-performing student, yields the lowest model performance in error identification. However, S6, also a high-performing student, achieves relatively better error identification results but receives the lowest LLM ratings for feedback quality. Meanwhile, S5 obtains the highest accuracy in correctness classification, though this may simply reflect that certain types of errors are more straightforward to detect, rather than indicating superior model capability. We hypothesize this disparity stems from problem-solving complexity. As shown in Figure 3, S2 employs the most equations and words per problem, which may account for the poor error identification performance. S6’s concise style with minimal verbal explanation may facilitate error localization but simultaneously limits the contextual information available for generating feedback. These findings suggest that different tasks may be influenced by different aspects of student problem-solving behavior.